# Language Models Encode Collaborative Signals in Recommendation

### v1

@@ C.1 Datasets
Books Movies & TV Video Games Industrial MovieLens-1M Book Crossing
@@
We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com [1] (i.e., Books, Movies & TV, Video Games, and Industrial), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing). Table 7 reports the data statistics of each dataset.
We divide the history interaction of each user into training, validation, and testing sets with a ratio of 4:3:3, and remove users with less than 20 interactions following previous studies [50]. We also remove items from the test and validation sets that do not appear in the training set, to address the cold start problem.
@@
Item Title Examples Books: Dismissed with Prejudice: A J.P. Beaumont Novel; Die for Love: A Jacqueline Kirby Novel of Suspense; The Cloud; Memories Before and After the Sound of Music: An Autobiography; Harry Potter and the Sorcerer’s Stone; Movies & TV: Batman Begins; Fantastic Four; Max Headroom: The Complete Series; Madagascar; Land of the Dead; King Kong; Video Games: Fighting Force; Tomb Raider II; Tomb Raider; WWF Warzone; Kartia: The Word of Fate; Snowboard Kids; Command & amp; Conquer: Tiberian Sun - PC; Final Fantasy VII; Grim Fandango - PC; Half-Life - PC; MovieLens-1M: Basquiat (1996); Tin Cup (1996); Godfather, The (1972); Supercop (1992); Manny & Lo (1996); Bound (1996); Carpool (1996); Book Crossing: Prague : A Novel; Chocolate Jesus; Wie Barney es sieht; To Kill a Mockingbird; Sturmzeit. Roman; A Soldier of the Great War; Pride and Prejudice (Dover Thrift Editions); Industrial: Jurassic Perisphinctes Ammonites from France; FS9140: Spinosaurus - Dinosaur Tooth 20-30mm; FS9410: USA Eocene, Fossil Fish (Knightia alt), A-grade; Delta 50-857 Charcoal Filter for 50-868; Hitachi RP30SA 7-1/2 Gallon Stainless Steel Industrial Shop Vacuum (Discontinued by Manufacturer); Makita 632002-4 14-Inch Cut-Off Wheels (5-Pack) (Discontinued by Manufacturer); PORTER-CABLE 740001801 4 1/2-Inch by 10yd 180 Grit Adhesive-Backed Sanding Roll;
@@
C.2 General Recommendation
C.2.1 Baselines
We incorporate a series of CF models as our baselines for general recommendation. These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec). For these LM-enhanced CF methods, we adopt the leading CF method XSimGCL as the backbone.
  • MF [55, 71] is the most basic CF model. It denotes users and items with ID-based embeddings and conducts matrix factorization with Bayesian personalized ranking (BPR) loss.
  • MultVAE [64] is a traditional CF model based on the variational autoencoder (VAE). It regards the item recommendation as a generative process from a multinomial distribution and uses variational inference to estimate parameters. We adopt the same model structure as suggested in the paper: 600 → 200 → 600.
  • LightGCN [56] is a light graph convolution network tailored for the recommendation, which deletes redundant feature transformation and activation function in NGCF [52].
  • SGL [83] introduces graph contrastive learning into recommender models for the first time. By employing node or edge dropout to generate augmented graph views and conduct contrastive learning between two views, SGL achieves better performance than LightGCN.
  • BC Loss [79] introduces a robust and model-agnostic contrastive loss, handling various data biases in recommendation, especially for popularity bias.
  • XSimGCL [70] directly generates augmented views by adding noise into the inner layer of LightGCN without graph augmentation. The simplicity of XSimGCL leads to a faster convergence speed and better performance.
  • KAR [48] enhances recommender models by integrating knowledge from large language models (LLMs). It generates textual descriptions of users and items and combine the LM representations with traditional recommenders using a hybrid-expert adaptor.
  • RLMRec [47] aligns semantic representations of users and items with the representations in CF models through a contrastive loss, as an additional loss trained together with the CF model. The fusion of semantic information and collaborative information brings performance improvement.
C.2.2 Ablation Study
@@
We conduct the same ablation study as introduced in Section 4.1 on Movies & TV and Video Games datasets. As illustrated in Figure 5, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.1.
C.2.3 The t-SNE Visualization Comparison
@@
Figure 6: The t-SNE visualization of representations on Movies & TV. (6(a)) The item representations in the LM space. (6(b)) The item representations obtained by replacing the MLP with a linear mapping matrix in AlphaRec. (6(c)) The item representations obtained from AlphaRec.
In this section, we aim to intuitively explore how the MLP in AlphaRec further helps in excavating collaborative signals in language representations, compared to the linear mapping matrix. We visualize the item representations from LMs, AlphaRec (w/o MLP), and AlphaRec in Figure 6, where AlphaRec (w/o MLP) denotes replacing the MLP with a linear mapping matrix. We observed that movies about superhero and monster cluster in all representation spaces, indicating both AlphaRec (w/o MLP) and AlphaRec capture the preference similarities between these items and preserve the clustering relationship. The difference between AlphaRec (w/o MLP) and AlphaRec lies in the ability to capture obscure preference similarities among items. As shown in Figure 6(a), homosexual movies are dispersed in the language space, indicating the possible semantic differences between them. AlphaRec successfully captures the preference similarities and gathers these items in the representation space, while AlphaRec (w/o MLP) remains some items dispersed. Moreover, AlphaRec outperforms AlphaRec (w/o MLP) by a large margin, as indicated in Figure 5(a). These results indicate that AlphaRec exhibits a more fine-grained preference capture ability with the help of nonlinear transformation.
C.3 Zero-shot Recommendation
C.3.1 Co-training on Multiple Datasets
@@
C.3.2 Baselines
Since previous works about zero-shot recommendation mostly focus on sequential recommendation [86, 85], we slightly modify two methods in sequential recommendation, ZESRec [38] and UniSRec [37] as our baselines. Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.
@@ C.3.2 Baselines
  • ZESRec [38] is the first work that defines the problem of zero-shot recommendation. To address this problem, this work introduces a hierarchical Bayesian model with representations from the pre-trained BERT.
  • UniSRec [37] aims to learn universal item representations from BERT, with parametric whitening and a MoE-enhanced adaptor. By pre-training on multiple source datasets, UniSRec can conduct zero-shot recommendation on various datasets in a transductive or inductive paradigm.
C.3.3 The Effect of Training Datasets
@@
Industrial MovieLens-1M Book Crossing
@@
C.4 User Intention Capture
C.4.1 Intention Query Generation
@@
The user intention query is a natural language sentence implying the target item of interest. For each item in the dataset, we generate a fixed user intention query. Following the previous work [40], we generate user intention queries with the help of ChatGPT [88]. As shown in Figure 7, we prompt ChatGPT in a Chain-of-Thought (CoT) [103] paradigm and adopt the output as the user intention query. We adopt a rule-based strategy to ensure the quality of generated queries, and regenerate the wrong query. Considering the huge amount of item title text, we use ChatGPT3.5 API for generating all queries for the budget’s sake.
C.4.2 Baseline
AlphaRec exhibits user intention capture abilities, although not specially designed for search tasks. We compare AlphaRec with TEM [89] which falls in the field of personalized search [87, 104].
  • TEM [89] uses a transformer to encode the intention query together with user history behaviors, which enables it to achieve better search results by considering the user’s historical interest.
C.4.3 Case Study
@@
C.4.4 Effect of the Intention Strength Alpha
@@
C.5 Trainig Cost
Table 10: Training cost of AlphaRec (seconds per epoch/in total).
Books Movies & TV Video Games Amazon-Mix
AlphaRec 40.1 / 1363.4 12.3 / 479.7 7.4 / 214.6 107.2 / 5788.8
Appendix D Hyperparameter Settings and Implementation Details
@@
We conduct all the experiments in PyTorch with a single NVIDIA RTX A5000 (24G) GPU and a 64 AMD EPYC 7543 32-Core Processor CPU. We optimize all methods with the Adam optimizer. For all ID-based CF methods, we set the layer numbers of graph propagation by default at 2, with the embedding size as 64 and the size of sampled negative items $|\mathcal{S}_{u}|$ as 256. We use the early stop strategy to avoid overfitting. We stop the training process if the Recall@20 metric on the validation set does not increase for 20 successive evaluations. In AlphaRec, the dimensions of the input and output in the two-layer MLP are 3072 and 64 respectively, with the hidden layer dimension as 1536. We apply the all-ranking strategy [105] for all experiments, which ranks all items except positive ones in the training set for each user. We search hyperparameters for baselines according to the suggestion in the literature. The hyperparameter search space is reported in Table 11. For these LM-enhanced models, KAR and RLMRec, we also search the hyperparameter of their backbone XSimGCL.
@@
Appendix E Broader Impact
The proposed AlphaRec can significantly improve the performance of zero-shot recommendation and the capability of user intent capture, offering a good approach to crafting more personalized recommendation results. One concern of AlphaRec is the potential for the representations generated by language models can be maliciously attacked, which may result in erroneous or unexpected recommendations. Therefore, we kindly advise researchers to cautiously check the quality of the language representations before using AlphaRec.

### v2

@@ C.1 Datasets
Books Movies & TV Video Games Industrial & Scientific MovieLens-1M Book Crossing
@@
We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com (Ni et al., 2019) (i.e., Books, Movies & TV, Video Games, and Industrial & Scientific), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing). Table 6 reports the dataset statistics.
We divide the history interaction of each user into training, validation, and testing sets with a ratio of 4:3:3, and remove users with less than 20 interactions following previous studies (Zhang et al., 2023b). We also remove items from the test and validation sets that do not appear in the training set, to address the cold start problem.
@@
Item Title Examples Books: Dismissed with Prejudice: A J.P. Beaumont Novel; Die for Love: A Jacqueline Kirby Novel of Suspense; The Cloud; Memories Before and After the Sound of Music: An Autobiography; Harry Potter and the Sorcerer’s Stone; Movies & TV: Batman Begins; Fantastic Four; Max Headroom: The Complete Series; Madagascar; Land of the Dead; King Kong; Video Games: USB Microphone for RockBand or Guitar Hero (PS3, Wii, Xbox360); Command &amp; Conquer: Tiberian Sun - PC; Tomb Raider III: Adventures of Lara Croft; Kartia: The Word of Fate; Snowboard Kids; Command & amp; Conquer: Tiberian Sun - PC; Final Fantasy VII; Grim Fandango - PC; Half-Life - PC; MovieLens-1M: Basquiat (1996); Tin Cup (1996); Godfather, The (1972); Supercop (1992); Manny & Lo (1996); Bound (1996); Carpool (1996); Book Crossing: Prague : A Novel; Chocolate Jesus; Wie Barney es sieht; To Kill a Mockingbird; Sturmzeit. Roman; A Soldier of the Great War; Pride and Prejudice (Dover Thrift Editions); Industrial & Scientific: Jurassic Perisphinctes Ammonites from France; FS9140: Spinosaurus - Dinosaur Tooth 20-30mm; FS9410: USA Eocene, Fossil Fish (Knightia alt), A-grade; Delta 50-857 Charcoal Filter for 50-868; Hitachi RP30SA 7-1/2 Gallon Stainless Steel Industrial Shop Vacuum (Discontinued by Manufacturer); Makita 632002-4 14-Inch Cut-Off Wheels (5-Pack) (Discontinued by Manufacturer); PORTER-CABLE 740001801 4 1/2-Inch by 10yd 180 Grit Adhesive-Backed Sanding Roll;
@@
C.2 Baselines
We incorporate a series of ID-based CF models as our baselines for general recommendation. These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec). For these LM-enhanced CF methods, we adopt the leading method XSimGCL as the backbone.
  • MF (Koren et al., 2009; Rendle et al., 2012) is the most basic CF model. It denotes users and items with ID-based embeddings and conducts matrix factorization with Bayesian personalized ranking (BPR) loss.
  • MultVAE (Liang et al., 2018) is a traditional CF model based on the variational autoencoder (VAE). It regards the item recommendation as a generative process from a multinomial distribution and uses variational inference to estimate parameters. We adopt the same model structure as suggested in the paper: 600 → 200 → 600.
  • LightGCN (He et al., 2021) is a light graph convolution network tailored for the recommendation, which deletes redundant feature transformation and activation function in NGCF (Wang et al., 2019b).
  • SGL (Wu et al., 2021) introduces graph contrastive learning into recommender models for the first time. By employing node or edge dropout to generate augmented graph views and conduct contrastive learning between two views, SGL achieves better performance than LightGCN.
  • BC Loss (Zhang et al., 2022) introduces a robust and model-agnostic contrastive loss, handling various data biases in recommendation, especially for popularity bias.
  • XSimGCL (Yu et al., 2024) directly generates augmented views by adding noise into the inner layer of LightGCN without graph augmentation. The simplicity of XSimGCL leads to a faster convergence speed and better performance.
  • KAR (Xi et al., 2023) enhances recommender models by integrating knowledge from LMs. It generates textual descriptions of users and items and combine the LM representations with traditional recommenders using a hybrid-expert adaptor.
  • RLMRec (Ren et al., 2024b) aligns semantic representations of users and items with the representations in CF models through a contrastive loss, as an additional loss trained together with the CF model. The fusion of semantic information and collaborative information brings performance improvement.
C.3 Ablation Study
@@
We conduct the same ablation study as introduced in Section 4 on Movies & TV and Video Games datasets. As illustrated in Figure 6, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.
C.4 The t-SNE Visualization Comparison
@@
Figure 7: The t-SNE visualization of representations on Movies & TV. (7(a)) The item representations in the LM space. (7(b)) The item representations obtained by replacing the MLP with a linear mapping matrix in AlphaRec. (7(c)) The item representations obtained from AlphaRec.
In this section, we aim to intuitively explore how the MLP in AlphaRec further helps in excavating collaborative signals in language representations, compared to the linear mapping matrix. We visualize the item representations from LMs, post-mapping representations from AlphaRec (w/o MLP), and post-mapping representations from AlphaRec in Figure 7, where AlphaRec (w/o MLP) denotes replacing the MLP with a linear mapping matrix. We observed that movies about superhero and monster cluster in all representation spaces, indicating both AlphaRec (w/o MLP) and AlphaRec capture the preference similarities between these items and preserve the clustering relationship. The difference between AlphaRec (w/o MLP) and AlphaRec lies in the ability to capture obscure preference similarities among items. As shown in Figure 7(a), homosexual movies are dispersed in the language space, indicating the possible semantic differences between them. AlphaRec successfully captures the preference similarities and gathers these items in the representation space, while AlphaRec (w/o MLP) remains some items dispersed. Moreover, AlphaRec outperforms AlphaRec (w/o MLP) by a large margin, as indicated in Figure 6(a). These results indicate that AlphaRec exhibits a more fine-grained preference capture ability with the help of nonlinear transformation.
Appendix D Exploring Potentials of Language Representations for Recommendation
D.1 Fast Convergence Speed
Table 7: Training cost of AlphaRec (seconds per epoch/in total).
Books Movies & TV Video Games Amazon-Mix
AlphaRec 40.1 / 1363.4 12.3 / 479.7 7.4 / 214.6 107.2 / 5788.8
D.2 Zero-shot Ability
D.2.1 Co-training on Multiple Datasets
@@
D.2.2 Baselines
Since previous works about zero-shot recommendation mostly focus on sequential recommendation (Kang & McAuley, 2018; Wang et al., 2019a), we slightly modify two methods in sequential recommendation, ZESRec (Ding et al., 2021) and UniSRec (Hou et al., 2022) as our baselines. Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.
@@ D.2.2 Baselines
  • ZESRec (Ding et al., 2021) is the first work that defines the problem of zero-shot recommendation. To address this problem, this work introduces a hierarchical Bayesian model with representations from the pre-trained BERT.
  • UniSRec (Hou et al., 2022) aims to learn universal item representations from BERT, with parametric whitening and a MoE-enhanced adaptor. By pre-training on multiple source datasets, UniSRec can conduct zero-shot recommendation on various datasets in a transductive or inductive paradigm.
D.2.3 Comparison with LLMRank
Table 8: Zero-shot performance comparison with LLMRank
MovieLens-1M Steam
NDCG@1 NDCG@5 NDCG@10 NDCG@20 NDCG@1 NDCG@5 NDCG@10 NDCG@20
LLMRank 0.2485 0.4115 0.5249 0.5612 0.3112 0.4413 0.5255 0.5302
AlphaRec 0.3919 0.6038 0.6543 0.6672 0.4450 0.6131 0.6394 0.6714
Imp .57.71% 46.73% 24.65% 18.89% 42.99% 38.93% 21.67% 26.63%
Table 8 illustrates the zero-shot recommendation performance compared with the LLM4Rec method LLMRank. We adopt the same setting of LLMRank, equipping 19 negative items for each positive item, and evaluate the NDCG on the candidate set. AlphaRec exhibits excellent zero-shot performance, significantly surpassing LLMRank. Moreover, the improvement over LLMRank exhibits a rising trend as the $K$ of NDCG decreases.
D.2.4 The Effect of Training Datasets
@@
Industrial & Scientific MovieLens-1M Book Crossing
@@
D.3 Intention-aware Ability
D.3.1 Intention Query Generation
@@
The user intention query is a natural language sentence implying the target item of interest. For each item in the dataset, we generate a fixed user intention query. Following the previous work (Hou et al., 2024a), we generate user intention queries with the help of ChatGPT (OpenAI, 2023). As shown in Figure 8, we prompt ChatGPT in a Chain-of-Thought (CoT) (Wei et al., 2022) paradigm and adopt the output as the user intention query. We adopt a rule-based strategy to ensure the quality of generated queries, and regenerate the wrong query. Considering the huge amount of item title text, we use ChatGPT3.5 API for generating all queries for the budget’s sake.
D.3.2 Baseline
AlphaRec exhibits user intention capture abilities, although not specially designed for search tasks. We compare AlphaRec with TEM (Bi et al., 2020) which falls in the field of personalized search (Ai et al., 2017; McAuley et al., 2015).
  • TEM (Bi et al., 2020) uses a transformer to encode the intention query together with user history behaviors, which enables it to achieve better search results by considering the user’s historical interest.
D.3.3 Case Study
@@
D.3.4 Effect of the Intention Strength Alpha
@@
Appendix E Hyperparameter Settings and Implementation Details
@@
We conduct all the experiments in PyTorch with a single NVIDIA RTX A5000 (24G) GPU and a 64 AMD EPYC 7543 32-Core Processor CPU. We optimize all methods with the Adam optimizer. For all ID-based CF methods, we set the layer numbers of graph propagation by default at 2, with the embedding size as 64 and the size of sampled negative items $|\mathcal{S}_{u}|$ as 256. We use the early stop strategy to avoid overfitting. We stop the training process if the Recall@20 metric on the validation set does not increase for 20 successive evaluations. In AlphaRec, the dimensions of the input and output in the two-layer MLP are 3072 and 64 respectively, with the hidden layer dimension as 1536. We apply the all-ranking strategy (Krichene & Rendle, 2020) for all experiments, which ranks all items except positive ones in the training set for each user. We search hyperparameters for baselines according to the suggestion in the literature. The hyperparameter search space is reported in Table 11. For these LM-enhanced models, KAR and RLMRec, we also search the hyperparameter of their backbone XSimGCL.